# references.bib

### v1

@@ V-C Implementation Details and Evaluation Settings
We measure precision/recall of observed predicates. Predicates that we consider are either unary or binary. In our implementation, for binary predicates the logic from Equation 6 directly applies, for unary predicates we set $v_{i}=v_{j}$ for uniformity with the $(v_{i},r,v_{j})$ edge definition. Instead of $b_{i}-b_{j}(\in\mathds{R}^{4})$, the spatial feature is then created through differences from coordinate to coordinate within the box, resulting in a feature vector $\in\mathds{R}^{6}$. For all one-shot experiments, to ensure there is a positive and negative reference, we sample a problem that contains each predicate type occurring in the test image at least once and is not always true.
To achieve the described object detector functionality, we use models from the GroundingDINO family. The class queries are passed to DINO-X [dinox] and the goal phrases to GroundingDINO 1.0. This is because DINO-X is stronger, but its API does not provide phrase grounding mode currently. For fair comparison with the baseline [vilain] that uses GroundingDINO for class queries, we upgrade its object detector to DINO-X, too.
For the LLM in the goal parser we use GPT-4o-mini which is sufficient for the task to convert the language instruction to PDDL format. Except for methods that are designed as zero-shot task planners or unless otherwise stated, all experiments are conducted in the one-shot setting, where one input-output example is provided. For the baselines, when we mention GPT-4o, we refer to gpt-4o-2024-11-20.
For the symbolic planner, we use Fast Downward [fastdownward], marked as “Solver”. As we are interested in evaluating the quality of state grounding, we do not allow re-prompting with error feedback from the planner to post-hoc alter state estimation.
@@ V-D1 PDDL State Generation
With respect to the problem setting, ViLaIn [vilain] is the most similar work to ours since their goal is to generate a PDDL problem consisting of objects, initial and goal state. Their methodological approach is, however, very different. It can be categorized under “reasoning with LLM informed by perception data”, requiring a GroundingDINO object detector, a BLIP2-captioner, and a GPT-4 reasoner. As shown in Table II(right), its success rates on task planning are $\leq 10\%$. Since they use the same planner as ours, the issue lies in its poor state grounding ability (middle). It suffers from three issues. First, the captioner is not conditioned on task and domain, producing irrelevant captions that do not allow for subsequent inference of the relevant predicates. Second, an LLM needs to write the state in PDDL, which can lead to violation of type rules or hallucinated or missing objects, documented in Figure 3 and problem validity scores $<1$ in Table III. Third, their decision to generate the goal state after initial state leads to error propagation from initial to goal state. All three problems are avoided by our state grounding framework design through (i) not relying on a captioner, (ii) writing initial state with DC-SGG instead of LLM, (iii) writing the goal state first.
@@ C B H C B H
Figure 3: Left: ViLaIn [vilain] detects false positive (FP) relation “knife-at-board” and misses (FN) relation “right_gripper-carry-knife”. Moreover, through its text-generative nature it can produce ungrounded hallucinations or forget already detected objects, here it forgets to include the “right_gripper”, dashed red box. Right: DC-SGGVQA with SOM labels to uniquely refer to instances while querying LMM.
@@ V-D2 LMM-As-Planner and State Estimator
VILA [vila] ask the model to first list task-relevant objects. Their reported prompt includes a rough description of constraints specific to their setups, so we need to adapt these sentences to our domain specification. Since we assume domain specification in PDDL format, but [vila] in natural language, we let an LLM generate a brief text description from the PDDL that matches the style of their prompt and verify that it is reasonable. We can only measure task planning, not state grounding, because their intermediate output does not contain triplet information. Results show that GPT-4o does not follow the domain-specific constraints for state grounding and planning, causing generation of invalid actions and leading to poor success rates $\leq 20\%$ on all domains. NeuroGround [neuroground] also uses one LMM for both state grounding and planning, but additionally assists plan generation with a symbolic engine similar to our solver. We evaluate their two variants LMM as planner vs. symbolic engine as planner. We observe that the symbolic engine can avoid LMM planning mistakes in the Blocksworld domain ($0.61$ vs. $0.37$) where the LLM often misinterprets the functionality of the four actions stack, unstack, pickup, putdown. However, at the same time parsing issues arise from the mismatch of state representation in weakly structured text modality and the highly structured language in which the parser and engine operate. An effect of this is that the LMM-planning succeeds in the cooking domain while its generated state cannot be parsed for the engine so that the engine variant scores lower here ($1.00$ vs. $0.88$).
VeriGraph [verigraph] operates, more similar to ours, strictly two-staged by first generating a scene graph and then planning on it. However, their planning is implemented as prompting the LMM for graph editing, and they need multiple iterations to correct errors through feedback from an additional validator function. As we do not allow these re-planning attempts, we only measure the state grounding accuracy. As with other GPT-4o-based methods, state grounding is accurate for the Cooking domain, while for Hanoi it fails to extract the correct number, color, and position of disks and pegs, leading to errors similar to those seen in other models like [vila, neuroground]. We present an example in Figure 4.
@@
Figure 4: The methods [vila, neuroground, verigraph] suffer from GPT-4o’s mistakes in determining the initial state for the test image (wrong predictions highlighted in red). [vila, neuroground] generate initial state as text representation, [verigraph] as scene graph through coding JSON, ours as domain-conditioned scene graph (bottom right), corresponding to a PDDL init state (bottom left).
@@ V-E Experiment on Predicate Classification through VQA
To provide more evidence that LMMs struggle with the required domain-specific, granular scene understanding, we show results for replacing our predicate classification through visual question answering (VQA) with an LMM, marked as DC-SGGVQA in Table II. Specifically, we ask a yes/no question for each possible grounded predicate. The initial state is then the set of predicates where the LMM answered yes. Since the questions are about specific instances that have been identified by the object detector, they must include references that can be uniquely associated with each instance. Therefore we use set-of-marks (SOM) [som] prompting for GPT-4o/MiniCPM-o-2_6 [minicpmo] and also try xGen-MM(BLIP3) [blip3] which was trained with interleaved bounding boxes. We show quantitative comparison in Table IV. With an example in Figure 3, state grounding results are better compared to ViLaIn for Blocksworld and Cooking, but the approach is still ineffective for two reasons. First, for predicates that take two object arguments, the set of grounded predicates grows quadratically w.r.t. objects. This forbids joint reasoning over the whole predicate set. Questions must be passed individually or batched, leading to the possibility of contradictory combinations like one disk being stacked on two pegs at the same time. Second, as shown in Figure 5, there are still various obvious wrong answers, suggesting a lack of elementary object-level perception ability. A series of other works [spatialvlm, llavagrounding, spatialrgpt] towards models with improved spatial and grounded question answering may help to eventually resolve this issue. Currently, however, with parameters below 8B, we found them to be inferior compared to Table IV.
@@ V-F Experiment on Open-Vocabulary SGG
In the above Section V-D and Section V-E we have observed that LMM-based methods face challenges as the number of objects and predicates in the scene increases. Classic scene graph generation methods do not have this problem as they natively operate on a large set of object and relationship proposals. Recent open-vocabulary scene graph generators like RLIPv2 [rlipv2] or Scene Graph-ViT [sgvit] are built similarly to open-vocabulary object detectors with additional relationship support, allowing to match object/relationship proposals with text queries. Open-vocab SGG would therefore constitute a more elegant solution to unify our scene parser in a single model. In DC-SGGOV we experiment with RLIPv2 [rlipv2] ParSeDA, Swin-L as zero-shot SGG as we identify no simple way to feed the one-shot example. The verb text queries are the predicate names within the domain specification and the object text queries are the same as for our object detector. Entity predictions are subjected to non-maximum suppression, and we set the score threshold for including a triplet score in the final predictions based on the empirically optimal value for each domain. Same as for our predicate classifier, we use the domain information to consider only triplets that may form a valid grounded predicate. Results show that in principle the solution is possible, but RLIPv2 struggles with the long-tail object types like the chopping knife in the Cooking domain or the peg in the Hanoi domain. Moreover, the meaning of predicate names might be ambiguous, e.g. one disk “on” another disk could mean directly on top or just somewhere above in the stack. In the results therefore no scene graph is entirely correct, preventing plan success. Nevertheless, we see efforts [scenespecificsg] to develop an interface for aligning open-vocab SGG models with the target concept as a promising direction to tackle the above issues.
@@ VI-A Extendability
3D. Our approach classifies predicates based on spatial features. In the dataset used, relevant relationships are distinguishable in 2D. However, in other environments, it may be necessary to incorporate 3D information to classify predicates. In this case, one can build spatial features from point clouds using a depth sensor, similar to the approach described in [spatialvlm]. Masks. Alternatively, in environments where fine-grained spatial information is more important, features could be constructed using segmentation masks, such as those produced by GroundedSAM, rather than relying solely on DINO-X bounding boxes. Goal Modality. In this paper we evaluated under instructions in text form, but our framework also supports goal images. The LLM-based goal parser is then replaced through a second scene parser from Section IV-A that determines the goal state through generating a scene graph of the goal image. Besides natural language and goal images, logical goal conditions as in [geometricsymbolic] can be directly provided, essentially skipping the goal parser. Closed Loop. Experiments in this paper were limited to an open-loop setting, but our framework can also run in a loop - more resource efficient than LMM approaches. Only the scene parser runs again to extract a new estimate of object set and initial state. The goal remains constant throughout execution, so the goal parser LLM only needs to be run once at the start.
@@ VI-B Limitations and Future Work
The evaluations in this paper were performed based on the dataset from the baseline [vilain], covering hand-crafted manipulation scenes. We acknowledge the quantitative results shown have limited implications on other task families and therefore in the next step, we want to evaluate in simulation, where we can (i) validate that the method works together with existing low-level skills and (ii) generate a dataset controlling for configurations to allow for more systematic analysis.

### v2

@@ V-C Implementation Details and Evaluation Settings
We measure precision/recall of observed predicates. Predicates that we consider are either unary or binary. In our implementation, for binary predicates the logic from Equation 6 directly applies, for unary predicates we set $v_{i}=v_{j}$ for uniformity with the $(v_{i},r,v_{j})$ edge definition. Instead of $b_{i}-b_{j}(\in\mathds{R}^{4})$, the spatial feature is then created through differences from coordinate to coordinate within the box, resulting in a feature vector $\in\mathds{R}^{6}$. For all one-shot experiments, to ensure there is a positive and negative reference, we sample a problem from ProDG-v that contains each predicate type occurring in the test image at least once and is not always true.
To achieve the described object detector functionality, we use models from the Grounding DINO family. The class queries are passed to DINO-X [24] and the goal phrases to Grounding DINO 1.0. This is because DINO-X is stronger, but its API does not provide phrase grounding mode currently. For fair comparison with the baseline [6] that uses GroundingDINO for class queries, we upgrade its object detector to DINO-X, too. For the LLM in the goal parser we use GPT-4o-mini which is sufficient for the task to convert the language instruction to PDDL format. Except for methods that are designed as zero-shot task planners or unless otherwise stated, all experiments are conducted in the one-shot setting, where one input-output example is provided. For the baselines, when we mention GPT-4o, we refer to gpt-4o-2024-11-20. For the symbolic planner, we use Fast Downward [23], marked as “Solver”. As we are interested in evaluating the quality of state grounding, we do not allow re-prompting with error feedback from the planner to post-hoc alter state estimation.
@@ V-D1 PDDL State Generation
With respect to the problem setting, ViLaIn [6] is the most similar work to ours since their goal is to generate a PDDL problem consisting of objects, initial and goal state. Their methodological approach is, however, very different. It can be categorized under “reasoning with LLM informed by perception data”, requiring a GroundingDINO object detector, a BLIP2-captioner, and a GPT-4 reasoner. As shown in Table II(right), its success rates on task planning are $\leq 10\%$. Since they use the same planner as ours, the issue lies in its poor state grounding ability (middle). It suffers from three issues. First, the captioner is not conditioned on task and domain, producing irrelevant captions that do not allow for subsequent inference of the relevant predicates. Second, an LLM needs to write the state in PDDL, which can lead to violation of type rules or hallucinated or missing objects, documented in Figure 3 and problem validity scores <1<1 in Table III. Third, their decision to generate the goal state after initial state leads to error propagation from initial to goal state. All three problems are avoided by our state grounding framework design through (i) not relying on a captioner, (ii) writing initial state with DC-SGG instead of LLM, (iii) writing the goal state first.
@@ C B H C B H
Figure 3: Left: ViLaIn [6] detects false positive (FP) relation “knife-at-board” and misses (FN) relation “right_gripper-carry-knife”. Moreover, through its text-generative nature it can produce ungrounded hallucinations or forget already detected objects, here it forgets to include the “right_gripper”, dashed red box. Right: DC-SGGVQA with SOM labels to uniquely refer to instances while querying LMM.
@@ V-D2 LMM-As-Planner and State Estimator
VILA [7] ask the model to first list task-relevant objects. Their reported prompt includes a rough description of constraints specific to their setups, so we need to adapt these sentences to our domain specification. Since we assume domain specification in PDDL format, but [7] in natural language, we let an LLM generate a brief text description from the PDDL that matches the style of their prompt and verify that it is reasonable. We can only measure task planning, not state grounding, because their intermediate output does not contain triplet information. Results show that GPT-4o does not follow the domain-specific constraints for state grounding and planning, causing generation of invalid actions and leading to poor success rates $\leq 20\%$ on all domains. NeuroGround [5] also uses one LMM for both state grounding and planning, but additionally assists plan generation with a symbolic engine similar to our solver. We evaluate their two variants LMM as planner vs. symbolic engine as planner. We observe that the symbolic engine can avoid LMM planning mistakes in the Blocksworld domain ($0.61$ vs. $0.37$) where the LLM often misinterprets the functionality of the four actions stack, unstack, pickup, putdown. However, at the same time parsing issues arise from the mismatch of state representation in weakly structured text modality and the highly structured language in which the parser and engine operate. An effect of this is that the LMM-planning succeeds in the cooking domain while its generated state cannot be parsed for the engine so that the engine variant scores lower here ($1.00$ vs. $0.88$).
VeriGraph [8] operates, more similar to ours, strictly two-staged by first generating a scene graph and then planning on it. However, their planning is implemented as prompting the LMM for graph editing, and they need multiple iterations to correct errors through feedback from an additional validator function. As we do not allow these re-planning attempts, we only measure the state grounding accuracy. As with other GPT-4o-based methods, state grounding is accurate for the Cooking domain, while for Hanoi it fails to extract the correct number, color, and position of disks and pegs, leading to errors similar to those seen in other models like [7, 5]. We present an example in Figure 4.
@@
Figure 4: The methods [7, 5, 8] suffer from GPT-4o’s mistakes in determining the initial state for the test image (wrong predictions highlighted in red). [7, 5] generate initial state as text representation, [8] as scene graph through coding JSON, ours as domain-conditioned scene graph (bottom right), corresponding to a PDDL init state (bottom left).
@@ V-E Experiment on Predicate Classification through VQA
To provide more evidence that LMMs struggle with the required domain-specific, granular scene understanding, we show results for replacing our predicate classification through visual question answering (VQA) with an LMM, marked as DC-SGGVQA in Table II. Specifically, we ask a yes/no question for each possible grounded predicate. The initial state is then the set of predicates where the LMM answered yes. Since the questions are about specific instances that have been identified by the object detector, they must include references that can be uniquely associated with each instance. Therefore we use set-of-marks (SOM) [25] prompting for GPT-4o/MiniCPM-o-2_6 [26] and also try xGen-MM(BLIP3) [27] which was trained with interleaved bounding boxes. We show quantitative comparison in Table IV. With an example in Figure 3, state grounding results are better compared to ViLaIn for Blocksworld and Cooking, but the approach is still ineffective for two reasons. First, for predicates that take two object arguments, the set of grounded predicates grows quadratically w.r.t. objects. This forbids joint reasoning over the whole predicate set. Questions must be passed individually or batched, leading to the possibility of contradictory combinations like one disk being stacked on two pegs at the same time. Second, as shown in Figure 5, there are still various obvious wrong answers, suggesting a lack of elementary object-level perception ability. A series of other works [28, 29, 30] towards models with improved spatial and grounded question answering may help to eventually resolve this issue. Currently, however, with parameters below 8B, we found them to be inferior compared to Table IV.
@@ V-F Experiment on Open-Vocabulary SGG
In the above Section V-D and Section V-E we have observed that LMM-based methods face challenges as the number of objects and predicates in the scene increases. Classic scene graph generation methods do not have this problem as they natively operate on a large set of object and relationship proposals. Recent open-vocabulary scene graph generators like RLIPv2 [31] or Scene Graph ViT [32] are built similarly to open-vocabulary object detectors with additional relationship support, allowing to match object/relationship proposals with text queries. Open-vocab SGG would therefore constitute a more elegant solution to unify our scene parser in a single model. In DC-SGGOV we experiment with RLIPv2 [31] ParSeDA, Swin-L as zero-shot SGG as we identify no simple way to feed the one-shot example. The verb text queries are the predicate names within the domain specification and the object text queries are the same as for our object detector. Entity predictions are subjected to non-maximum suppression, and we set the score threshold for including a triplet score in the final predictions based on the empirically optimal value for each domain. Same as for our predicate classifier, we use the domain information to consider only triplets that may form a valid grounded predicate. State grounding results in Table II show that in principle the solution is possible. However, no scene graph is entirely correct, which would be a necessity for plan success, leading to the reported zero sucess rates. RLIPv2 struggles with the long-tail object types like the chopping knife in the Cooking domain or the peg in the Hanoi domain. Moreover, the meaning of predicate names might be ambiguous, e.g. one disk “on” another disk could mean directly on top or merely somewhere above in the stack. We see efforts [33] to develop an interface for aligning object-centric open-vocab models with the target concept as a promising direction to tackle this issue.
@@ VI-A Extendability
We briefly explore the potential for extended variants.
3D. Our approach classifies predicates based on spatial features. In the dataset used, relevant relationships are distinguishable in 2D. However, in other environments, it may be necessary to incorporate 3D information to classify predicates. In this case, one can build spatial features from point clouds using a depth sensor, similar to the approach described in [28]. Masks. Alternatively, in environments where fine-grained spatial information is more important, features could be constructed using segmentation masks, such as those produced by GroundedSAM, rather than relying solely on DINO-X bounding boxes. Goal Modality. In this paper we evaluated under instructions in text form, but our framework also supports goal images. The LLM-based goal parser is then replaced through a second scene parser from Section IV-A that determines the goal state through generating a scene graph of the goal image. Besides natural language and goal images, logical goal conditions as in [14] can be directly provided, essentially skipping the goal parser. Closed Loop. Experiments in this paper were limited to an open-loop setting, but our framework can also run in a loop - more resource efficient than LMM approaches. Only the scene parser runs again to extract a new estimate of object set and initial state. The goal remains constant throughout execution, so the goal parser LLM only needs to be run once at the start.
@@ VI-B Limitations and Future Work
The evaluations in this paper were performed based on the dataset from the baseline [6], covering hand-crafted manipulation scenes. We acknowledge the quantitative results shown have limited implications on other task families and therefore in the next step, we want to evaluate in simulation, where we can (i) validate that the method works together with existing low-level skills and (ii) generate a dataset controlling for configurations to allow for more systematic analysis.